A stone is projected vertically upwards from the ground and 3s later another stone is projected with same speed.The 2 meet at a height 20m above the ground.Find the velocity of projection.

Grade:11

1 Answers

Jaya IITK
12 years ago

displacement same

s1=s2=20

ut1-1/2g t1^2=20

3u-1/2 gt^2=20...... eqn. 1

 

s2=20

 

u(t-3)-1/2 g (t-3)^2=20 ...... eqn. 2 

 

 

solving we get u

 

Plzzzz approve my answer by clicking on yes below!!!!!!!!!!!!